OpenStudy (anonymous):

Show that the function f(x) is a growth function for x<5 . f(x)=5-x/5^-x

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(x)=\frac{5-x}{5^{-x}} }\) this?

OpenStudy (anonymous):

yes sir

OpenStudy (solomonzelman):

OK. I would assume that "growth function for t>5" means that the function is increasing for (all) t>5.

OpenStudy (solomonzelman):

We can actually legitimately prove this using calculus, if this is a calculus question. Is it calculus?

OpenStudy (anonymous):

no sir algebra 2

OpenStudy (solomonzelman):

Alright.

OpenStudy (solomonzelman):

Well, then we can't "prove" the hypothesis. However, you could show that the function is going to be positive, and the amount will be always increasing for all x>5.

OpenStudy (solomonzelman):

I will first re-write the function as, \(\color{#000000}{ \displaystyle f(x)=(5-x)5^{x} }\)

OpenStudy (solomonzelman):

Do you understand how I was able to write the f(x) this way?

OpenStudy (anonymous):

no sir

OpenStudy (solomonzelman):

There is a rule: \(\color{#000000}{ \displaystyle a^{-b}=\frac{1}{a^b} }\) So, here I would get: \(\color{#000000}{ \displaystyle f(x)=\frac{5-x}{5^{-x}} }\) \(\color{#000000}{ \displaystyle f(x)=(5-x)\times \frac{1}{5^{-x}} }\) \(\color{#000000}{ \displaystyle f(x)=(5-x)\times \frac{1}{\frac{1}{5^x}} }\) \(\color{#000000}{ \displaystyle f(x)=(5-x)\times 5^x }\)

OpenStudy (solomonzelman):

How about now?

OpenStudy (anonymous):

yes sir i understand

OpenStudy (solomonzelman):

Ok, so your function, really is: \(\color{#000000}{ \displaystyle f(x)=(5-x)5^x }\) And you are asked to show that when x <5, the function is increasing.

OpenStudy (anonymous):

all it says is to show that the fuction f(x) is a growth function for x<5

OpenStudy (solomonzelman):

Yes, and I am interpreting that wording, as showing that the function is increasing.

OpenStudy (anonymous):

oh ok continue sir

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle f(x)=(5-x)5^x }\) I can do this by plugging in some values. (This is not really a proof, because the real proof is calculus, however it is a way of showing the hypothesis is true, so let's begin ...) \(\color{#000000}{ \displaystyle f(\color{red}{-3})=(5-(\color{red}{-3}))5^{\color{red}{-3}} =8\cdot \frac{1}{5^3}=\frac{8}{125} }\) \(\color{#000000}{ \displaystyle f(\color{red}{-2})=(5-(\color{red}{-2}))5^{\color{red}{-2}} =7\cdot \frac{1}{5^2}=\frac{7}{25} }\) and so on ...

OpenStudy (solomonzelman):

Plug 4 or 5 values into f(x) (that are smaller than 5), and you will see that the smaller the input the smaller the output (and the bigger the input, the bigger the output.) *** (((Note that for example, -6 is less than -3, and -2 is less than 1 .... and that should precisely clarify what I mean by "bigger" and "smaller")))

OpenStudy (solomonzelman):

when you get --> the bigger the input the bigger the output, you may than conclude that the function (for x<5) increases (or, "a growth function")

OpenStudy (solomonzelman):

Also, if you want a similar example, let me know.

OpenStudy (anonymous):

i have another question i will post it will you help me btw god bless you and thanks

OpenStudy (solomonzelman):

Not a problem. If I have time, I'll take a look:)